OpenStudy (anonymous):

What is the simplified version of sqrt(-80)/sqrt(10)? Does it include i?

OpenStudy (espex):

Yes it does.

OpenStudy (anonymous):

You can first take out the -1 out like so \(\sqrt{-80}\sqrt{10} = \sqrt{-1}\sqrt{80}\sqrt{10}\) then you can further simplify it. also, \(\sqrt{-1} = i\), so you can continue from there.

OpenStudy (anonymous):

Okay, so it's like \[i \sqrt{80}/\sqrt{10}\] But then what do I do? Oh, do I bring it to \[i4\sqrt{5}/\sqrt{10}\]?

OpenStudy (anonymous):

\[\frac{80}{10}=8\]

OpenStudy (anonymous):

so you have \[\sqrt{8}i\] or if you prefer \[2\sqrt{2}i\]

OpenStudy (anonymous):

Oh, that's it? I see! Thank you!

OpenStudy (anonymous):

yw